# FAST: An Efficient Scheduler for All-to-All GPU Communication

### v2

@@ 5 Evaluation
Testbed. We deploy FAST on two clusters with GPU Direct RDMA [4]: (i) NVIDIA cluster: 4 servers with NVIDIA H200 GPUs [38], each with 8 GPUs and 8 NICs, connected via 400 Gbps InfiniBand with credit-based flow control [35] and 4 KB MTU. Intra-server scale-up uses NVLink via NVSwitch, with a 9:1 scale-up to scale-out bandwidth ratio (450 GBps vs. 50 GBps). The scheduler runs on Intel Xeon Platinum 8468 CPUs. (ii) AMD cluster: 4 servers with AMD MI300X GPUs [2], each with 8 GPUs and 8 ConnectX-6 NICs, connected via 100 Gbps RoCEv2 Ethernet with out-of-the-box DCQCN as congestion control and 1 KB MTU. Intra-server scale-up is a fully connected Infinity Fabric mesh, with a 35:1 bandwidth ratio (448 GBps vs. 12.5 GBps). The scheduler runs on AMD EPYC 9534 CPUs.
Libraries & dependencies. We provide a unified Python API, all_to_all_FAST, mirroring PyTorch’s all_to_all_single for seamless integration into existing models. Implementation details differ by hardware: (i) On H200, scale-up uses CUDA IPC [34] and scale-out uses NVSHMEM [39]. (ii) On MI300X, both scale-up and scale-out are implemented with RCCL [5]. The transfer pipeline is realized with multiple CUDA/HIP streams and careful cross-stream synchronization.
Integration into MoE systems. FAST operates in a distributed fashion due to computational determinism: given the same traffic matrix, each GPU’s control thread independently computes the identical global schedule, removing the need for a central coordinator. Only the traffic matrix (a compact integer array) must be synchronized—not the schedule itself.
This integration is seamless in Megatron-LM [45], where GPUs already construct the traffic matrix before every alltoallv. FAST simply consumes this existing matrix, does the scheduling, and executes the transfers.
Workloads. We evaluate FAST on both synthetic and real MoE workloads. For synthetic workloads, we model skewness by varying GPU-pair transfer sizes using two distributions: (i) random alltoallv with uniformly-distributed sizes, and (ii) skewed alltoallv with Zipfian-distributed sizes. For real workloads, we focus on MoE training, where prior work identifies alltoallv as the primary bottleneck [20, 25, 23], and report the end-to-end throughput improvements achieved by FAST.
@@
Metrics. Our primary metric is algorithmic bandwidth, widely used in prior work [3, 44]. It captures how fast a transfer completes, defined as . Because skewed alltoallv may have variable per-GPU volumes, this average normalizes across all GPUs. Algorithmic bandwidth can exceed the raw scale-out link bandwidth, since part of the transfer completes locally over the faster scale-up fabric. For example, in a 4-node cluster with 50 GBps scale-out links, if 25% of the traffic is intra-server, the optimal algorithmic bandwidth is GBps.
@@
Baselines. We compare FAST against two classes of baselines: (i) Solver-based schedulers: TACCL [44] and TE-CCL [27], which used constraint-based solvers for scheduling. We evaluate them on both NVIDIA and AMD clusters. (ii) Industry libraries: On NVIDIA: NCCL [3] (recent version 2.27.3), DeepEP [46] (from DeepSeek [14]), and MSCCL [13]. On AMD: RCCL [5], SpreadOut [42] (abbreviated as ‘SPO’), and MSCCL.
Since DeepEP is NVIDIA-only, we cannot test it on AMD. NCCL is highly optimized and outperforms SpreadOut, so we omit SpreadOut results on NVIDIA. By contrast, RCCL’s alltoallv on AMD lacks scheduling (all flows launch concurrently), making SpreadOut a stronger baseline that we include explicitly.
@@ 5.1.1 Performance Under Different Transfer Sizes
FAST consistently achieves the best alltoallv performance on both NVIDIA and AMD testbeds. For baselines, NCCL, DeepEP, SpreadOut, and RCCL support alltoallv natively, while TACCL, TE-CCL, and MSCCL only support balanced All-to-All. Adapting the solvers to skewed alltoallv can be done in two ways: (i) explicitly encoding variable flow sizes into the problem formulation, or (ii) padding all flows to a uniform size so the solver sees a balanced workload (padding data is used only for scheduling, not for actual transfers). We attempted the first approach, but it made the solvers—already slow on balanced workloads (e.g., TACCL needs over 30 minutes for 32 GPUs)—incapable of finishing within a reasonable time. Thus, we adopt the padding method to simplify workloads for solver-based schedulers.
@@
We evaluate with per-GPU message sizes from 100 MB to 1 GB, representative of typical workloads reported by prior work [25, 46]. The improvement factors vary across testbeds, reflecting differences in both network hardware and software implementation, as discussed below.
@@
Result on NVIDIA testbed. As shown in Figure˜13(a), FAST achieves the highest algorithmic bandwidth under random workloads. It slightly outperforms NCCL by 1.01–1.1, and exceeds DeepEP (1.0–1.4) and TACCL (1.5–1.7). Performance generally improves with larger transfers, as scale-out links saturate more easily and staging overheads—especially in FAST—are amortized.
NCCL with PXN [28] is highly optimized: data is aggregated at the source server before being forwarded across scale-out links. While NCCL does not explicitly rebalance like FAST, PXN aggregation smooths skew in lightly skewed workloads. As a result, NCCL comes close to FAST under this workload despite lacking explicit skew handling.
DeepEP offers separate implementations for dispatch and combine alltoallv, both with similar performance. However, its design routes all tokens through proxy GPUs at the source server via scale-up, which under skew creates hotspots on certain proxies and their scale-up links, reducing throughput (as confirmed by DeepEP’s own NVLink profiling).
Solver-based schedulers (TACCL, TE-CCL) reconstruct skewed alltoallv into balanced All-to-All via padding. Although this allows near-optimal schedules in theory, padded traffic does not contribute to real transfers and delays actual data. In practice, their schedules achieve only about half of FAST’s throughput.
@@
Straggler effects intensify under skew. With Zipfian workloads (Figure˜13(b)), FAST outperforms NCCL by 1.2–1.3, DeepEP by 1.15–1.4, and TACCL by over 3. The gap with NCCL widens under medium to high skew: even after PXN aggregation, residual imbalance produces stragglers that limit NCCL’s efficiency.
@@
Results on AMD Testbed. FAST again achieves the best performance under random workloads (Figure˜14(a)), surpassing TACCL by 1.3–1.8, TE-CCL by 1.6–2.3, SpreadOut by 1.9–2.1, and RCCL by 1.1–10. As with NVIDIA, most algorithms benefit from larger transfers.
@@
Under skewed workloads (Figure˜14(b)), FAST extends its lead, outperforming TACCL by 2.9–3.8, TE-CCL by 3.6–4.7, SpreadOut by 2.5–2.8, and RCCL by 1.3–2.6. Interestingly, RCCL performs relatively better here than under random workloads: skew concentrates traffic into a few elephant flows while leaving most as short mice transfers, reducing widespread collisions and easing incast pressure.
@@ 5.1.3 Performance under Different Skewness
On the AMD testbed, we compare FAST with TACCL, SpreadOut, and RCCL across different skewness levels (Figure˜15(a)). FAST consistently delivers the best performance, outperforming RCCL by 1.6–10, SpreadOut by 2.1–3.1, and TACCL by 2.1–4.5 (TE-CCL omitted here as it performs slightly worse than TACCL).
The performance gap reflects how each system handles stragglers: (i) For FAST, stronger skew increases the data that must be rebalanced, lengthening the balancing phase. As shown in the breakdown of Figure˜15(b), balancing time grows with skewness but remains modest: even at factor 0.9, balancing plus redistribution accounts for under 8% of scale-out time (and under 5% in most cases). Because scale-out dominates and runs at full efficiency, FAST stays within 1.05 of the optimal time on this AMD testbed. (ii) TACCL relies on padding, so heavier skew requires more padding, lowering effective efficiency. (iii) SpreadOut suffers as skew enlarges disparity across flows within each stage, causing stage time—and thus overall transfer time—to be dominated by stragglers. (iv) RCCL exhibits the opposite trend: higher skew generates more mice flows, which are easily absorbed by switch buffers. This reduces incast severity on the AMD testbed, improving RCCL’s performance.
@@ 5.2 End-to-End Performance
To evaluate FAST in an end-to-end setting, we integrate it into Megatron-LM [45] on the AMD testbed to perform on-the-fly scheduling for every alltoallv communication during MoE training. We compare against PyTorch’s [40] default all_to_all_single operator, which uses RCCL as the backend. Solver-based approaches cannot be integrated due to their prohibitive scheduling overhead.
We vary two key MoE configurations to study how FAST behaves under different training scenarios: (i) Expert parallelism (EP): We sweep EP from 16 to 24 to 32, which directly determines the scale of alltoallv. Under the configuration where each GPU hosts one expert (like DeepSeek [14]), this corresponds to scaling the transfer from 16 GPUs (2 servers) to 32 GPUs (4 servers). (ii) Top-K routing: In MoE, each input token is routed to the Top-K most relevant experts; larger increases token replication and thus flow size in the alltoallv workload. For this experiment, we fix EP at 32 and vary . Other workload-related parameters, such as batch size, sequence length, and data type, similarly affect communication size and thus training performance.
As shown in Figure˜16(a), FAST delivers a 1.18–4.48 speedup in end-to-end training throughput (under Top-2 routing) across different EP levels. Two main trends emerge: (i) Training throughput (left y-axis) decreases as EP increases. This is expected since higher EP involves more GPUs, more servers, and thus more scale-out traffic, which lowers communication efficiency and increases GPU idle time in Megatron-LM. (ii) The baseline degrades sharply as EP grows due to escalating incast. RCCL performs no scheduling across transfers, leaving congestion entirely to the network layer. For example, with EP16, a receiver GPU/NIC handles up to 8 concurrent flows, while with EP32 this rises to 24. With only out-of-the-box DCQCN as congestion control, this causes severe throughput collapse.
As shown in Figure˜16(b), FAST outperforms the baseline by 1.75–7.88 across different Top-K routing policies. Notably, FAST and RCCL follow opposite trends: (i) With FAST, increasing enlarges flow sizes in the alltoallv workload, which helps amortize staging overhead and more readily saturates link bandwidth, thereby boosting throughput (consistent with Figure˜14). (ii) With RCCL, however, larger exacerbates flow collisions and congestion, leading to further degradation in communication efficiency.
@@ 5.3 Scheduling Overhead
Scheduling runtime. As shown in Figure˜17, FAST scales to 320 GPUs with only 77 ms of overhead—faster than SyCCL, the fastest solver-based scheduler, which already takes 3.6s at just 16 GPUs. In contrast, earlier solver-based methods generally fail to scale beyond 64 GPUs, rendering them unusable for moderate expert-parallelism levels such as EP96 or EP128 [6].
@@
Memory overhead. FAST also requires additional memory for temporary buffers that hold rebalanced or redistributed data. Under random workloads, this overhead is about of the original alltoallv buffer size. In practice, the impact is minimal: typical alltoallv buffers are under 1 GB (e.g., in DeepEP [46]), so the extra cost is less than 300 MB. On modern GPUs such as the NVIDIA H200 [38] with 141 GB of memory, this represents under 0.22% of total capacity—an acceptable tradeoff for the performance gains.
5.4 Simulation at Larger Scales and Variable Bandwidths
@@
We use simulation to evaluate FAST beyond the limits of our physical testbeds, exploring both larger scales and different scale-up/scale-out bandwidth configurations. The simulator follows the analytical framework widely used in prior work such as TE-CCL and TACCL [44, 27]: given a schedule with a sequence of transfer steps (each with a defined size), the completion time is computed by summing per-step costs. Each cost consists of a fixed link wake-up delay plus the transmission time ().
@@
We focus on scenarios that are infeasible for solver-based schedulers, which fail to scale. Accordingly, we compare FAST against SpreadOut and an optimal bandwidth bound. This bound assumes infinitely fast scale-up links, so any intra-server transfers complete instantly. The remaining bottleneck is solely scale-out, with optimal time defined as the maximum sender or receiver load (after balancing) divided by the scale-out bandwidth.
@@
Performance at Larger Scale. We first scale the number of GPUs in alltoallv, with each GPU pair transmitting 50 MB on average in a random workload, simulated on a 400 Gbps scale-out network and a 450 GBps scale-up network (H200). As shown in Figure˜18(a), FAST stays within 5% of optimal when scheduling time is excluded (“FAST raw”). Including scheduling time, the gap widens to 10% at larger scales, since scheduling cost grows faster than workload completion time (which increases only linearly with GPU count). We leave on-the-fly scheduling at extreme scale as future work. By contrast, SpreadOut achieves only about half of FAST ’s throughput.
Overall performance declines as scale grows, driven by the rising share of traffic traversing the slower scale-out network.
@@
Performance under Varying Scale-up/Scale-out Ratios. We also evaluate FAST under different scale-up/scale-out bandwidth ratios on a 32-GPU setup. As shown in Figure˜18(b), normalized bandwidth is reported relative to scale-out capacity; values can exceed 1 since about 25% of traffic is intra-server, giving an upper bound near 1.25. Performance improves as the scale-up/scale-out ratio increases, as faster scale-up links further reduce balancing and redistribution overhead. These results indicate that FAST sustains high efficiency on modern GPUs with advanced scale-up interconnects.

### v3

@@ 5 Evaluation
Testbed. (i) NVIDIA cluster: 4 servers with NVIDIA H200 GPUs [h200], each with 8 GPUs, interconnected by 400 Gbps InfiniBand with credit-based flow control [infinibandflowcontrol] and 4 KB MTU. Intra-server scale-up uses NVLink, with a 9:1 scale-up to scale-out bandwidth ratio (450 GBps vs. 50 GBps). The scheduler runs on Intel Xeon Platinum 8468 CPUs. (ii) AMD cluster: 4 servers with AMD MI300X GPUs [mi300x], each with 8 GPUs, connected via 100 Gbps RoCEv2 Ethernet with out-of-the-box DCQCN [dcqcn] as congestion control and 1 KB MTU. Intra-server scale-up is a fully connected Infinity Fabric mesh, with a 35:1 bandwidth ratio (448 GBps vs. 12.5 GBps). The scheduler runs on AMD EPYC 9534 CPUs. Each GPU has its dedicated NIC with GPU Direct RDMA [gpudirect] in both clusters.
Libraries & dependencies. We provide a Python API, all_to_all_FAST, mirroring PyTorch’s all_to_all_single for integration into existing models. Implementations of data transfers differ by hardware: (i) On H200, scale-up/scale-out uses CUDA IPC [cuda]/NVSHMEM [nvshmem] respectively. (ii) On MI300X, both use RCCL [rccl]. The transfer pipeline is implemented using multiple CUDA streams with synchronizations.
Integration into MoE systems. FAST operates in a distributed fashion: given the same traffic matrix, each GPU independently computes the identical global schedule, eliminating the need for a central coordinator. Only the traffic matrix—a compact integer array—must be synchronized; the schedule itself does not need to be exchanged.
This integration is natural in MoE frameworks such as Megatron-LM [megatron-lm], which already materialize the per-alltoallv traffic matrix before each dispatch. Specifically, Megatron-LM performs an All-Gather of per-expert token counts (e.g., num_global_tokens_per_expert [megatron-lm-allgather]), from which the full traffic matrix can be constructed. Importantly, this All-Gather is not introduced by FAST. It is already required by the baseline NCCL alltoallv implementation to compute receive counts and buffer offsets, since each GPU knows how many tokens it sends but not how many it will receive from others. FAST simply consumes this existing traffic matrix, synthesizes a schedule, and executes the transfers.
Workloads. We evaluate FAST on both synthetic and real MoE workloads. For synthetic workloads, we model skewness by varying GPU-pair transfer sizes using two distributions: (i) random alltoallv with uniformly-distributed sizes, and (ii) skewed alltoallv with Zipfian-distributed sizes. For real workloads, we focus on MoE training, where prior work identifies alltoallv as the primary bottleneck [tutel, mixnet, gshard], and report the end-to-end throughput improvements.
@@
Metrics. Our primary metric is algorithmic bandwidth, widely used in prior work [nccl, taccl]. It captures how fast a transfer completes, defined as . Because skewed alltoallv may have variable per-GPU volumes, this average normalizes across all GPUs. Algorithmic bandwidth can exceed the raw scale-out link bandwidth, since part of the transfer completes locally over the faster scale-up fabric. For example, in a 4-node cluster with 50 GBps scale-out links, if 25% of the traffic is intra-server, the optimal algorithmic bandwidth is GBps. Higher is better.
@@
Baselines. We compare FAST against two classes of baselines: (i) Solver-based schedulers: TACCL [taccl] and TE-CCL [teccl], which used constraint-based solvers for scheduling. We evaluate them on both NVIDIA and AMD clusters. (ii) Industry libraries: On NVIDIA: NCCL [nccl] (version 2.27.3), DeepEP [deepep] (from DeepSeek [deepseek]), and MSCCL [msccl]. On AMD: RCCL [rccl], SpreadOut [mpi] (abbreviated as ‘SPO’), and MSCCL. Since DeepEP is NVIDIA-only, we do not evaluate it on AMD. NCCL outperforms SpreadOut on NVIDIA, so we omit SpreadOut there; on AMD, RCCL’s alltoallv lacks effective scheduling, making SpreadOut a stronger baseline that we include.
@@ 5.1.1 Performance Under Different Transfer Sizes
FAST consistently achieves the best alltoallv performance on both NVIDIA and AMD testbeds. For baselines, NCCL, DeepEP, SpreadOut, and RCCL support alltoallv natively, while TACCL, TE-CCL, SyCCL, and MSCCL only support balanced All-to-All. Adapting the solvers to skewed alltoallv can be done in two ways: (i) explicitly encoding variable flow sizes into the problem formulation, or (ii) padding all flows to a uniform size so the solver sees a balanced workload (padding data is used only for scheduling, not for actual transfers). We attempted the first approach, but it made the solvers—already slow on balanced workloads (e.g., TACCL needs over 30 minutes for 32 GPUs)—incapable of finishing within a reasonable time. Thus, we adopt padding to simplify workloads for solver-based schedulers.
@@
We evaluate with per-GPU message sizes from 100 MB to 1 GB, representative of typical workloads reported by prior work [mixnet, deepep]. The improvement factors vary across testbeds, reflecting differences in both network hardware and software implementation, as discussed below.
@@
Result on NVIDIA testbed. As shown in Figure˜12(a), FAST achieves the best algorithmic bandwidth under random workloads. It slightly outperforms NCCL by 1.01–1.1, and exceeds DeepEP (1.5–1.9) and TACCL (1.5–1.7). Performance improves with larger transfers, as scale-out links saturate more easily and staging overheads in FAST are amortized.
NCCL with PXN [ncclpxn] employs sender-side aggregation, consolidating outgoing flows at proxy GPUs before traversing scale-out links. By aggregating traffic across flows, PXN reduces per-GPU variance and mitigates mild skew. As a result, under mildly skewed workloads, NCCL can approach FAST’s performance even without explicit traffic rebalancing.
DeepEP [deepep] places aggregation and fan-out on the receiver side. Data are first delivered to ingress GPUs on the destination server and then forwarded via NVLink to their target GPUs. Under skew, multiple ingress GPUs may concurrently forward large volumes to the same targets, causing NVLink receive contention and local hotspots that limit throughput—as observed by DeepEP’s own NVLink runtime profiler.
Solver-based schedulers (e.g., TACCL, TE-CCL) convert skewed alltoallv into a fictitiously balanced All-to-All via padding, which significantly reduces synthesis time and enables practical schedule generation. However, the padded transfers do not correspond to real data movement and still occupy communication slots, delaying actual transfers. So their performances are much lower than FAST in practice.
@@
Straggler effects intensify under skew. Under Zipfian workloads (Figure˜12(b)), FAST outperforms NCCL by 1.2–1.3, DeepEP by 1.2–1.5, and TACCL by over 3. The performance gap with NCCL widens as the workload shifts from uniform to Zipfian: even with PXN aggregation, residual imbalance introduces stragglers that limit NCCL’s efficiency.
@@
Results on AMD testbed. FAST again achieves the best performance under random workloads (Figure˜13(a)), surpassing TACCL by 1.3–1.8, TE-CCL by 1.6–2.3, SpreadOut by 1.9–2.1, and RCCL by 1.1–10. As with NVIDIA, most algorithms benefit from larger transfers.
@@
Under skewed workloads (Figure˜13(b)), FAST extends its lead, outperforming TACCL by 2.9–3.8, TE-CCL by 3.6–4.7, SpreadOut by 2.5–2.8, and RCCL by 1.3–2.6. Interestingly, RCCL performs relatively better here than under random workloads: skew concentrates traffic into a few elephant flows while leaving most as short mice transfers, reducing widespread collisions and easing incast pressure.
@@ 5.1.3 Performance under Different Skewness
On the AMD testbed, we compare FAST with TACCL, SpreadOut, and RCCL across different skewness levels (Figure˜14(a)). FAST consistently delivers the best performance, outperforming RCCL by 1.6–10, SpreadOut by 2.1–3.1, and TACCL by 2.1–4.5 (TE-CCL omitted here as it performs slightly worse than TACCL).
The performance gap reflects how each system handles stragglers. (i) For FAST, increased skew lengthens the balancing phase, but the overhead remains modest as shown in Figure˜14(b): even at skew factor 0.9, balancing and redistribution account for under 8% of scale-out time (under 5% in most cases). Since scale-out dominates and runs at full efficiency, FAST remains within 1.08 of optimal. (ii) TACCL degrades under skew because heavier skew requires more padding, reducing effective efficiency. (iii) SpreadOut suffers as skew amplifies per-stage imbalance, causing overall transfer time to be dominated by stragglers. (iv) RCCL shows the opposite trend: higher skew produces more mice flows, whose contention is absorbed by switch buffers, reducing incast severity and improving performance on the AMD testbed.
@@ 5.2 End-to-End Performance
To evaluate FAST in an end-to-end setting, we integrate it into Megatron-LM [megatron-lm] on the AMD testbed to perform on-the-fly scheduling for every alltoallv communication during MoE training. We compare against PyTorch’s [pytorch] default all_to_all_single operator, which uses RCCL as the backend. Solver-based approaches cannot be integrated due to their prohibitive scheduling overhead.
We vary two key MoE configurations to study how FAST behaves under different training scenarios: (i) Expert parallelism (EP): We sweep EP from 16 to 24 to 32, which directly determines the scale of alltoallv. Under the configuration where each GPU hosts one expert (like DeepSeek [deepseek]), this corresponds to scaling the transfer from 16 GPUs (2 servers) to 32 GPUs (4 servers). (ii) Top-K routing: In MoE, each input token is routed to the Top-K most relevant experts; larger increases token replication and thus flow size in the alltoallv workload. For this experiment, we fix EP at 32 and vary . Other workload-related parameters, such as batch size and sequence length, similarly affect communication size and training performance.
As shown in Figure˜15(a), FAST delivers a 1.18–4.48 speedup in end-to-end training throughput (under Top-2 routing) across different EP levels. Two main trends emerge: (i) Training throughput (left y-axis) decreases as EP increases. This is expected since higher EP involves more GPUs, more servers, and thus more scale-out traffic, which lowers communication efficiency and increases GPU idle time. (ii) The baseline degrades sharply as EP grows due to escalating incast. RCCL performs no scheduling across transfers, leaving congestion entirely to the transport layer. For example, with EP16, a receiver GPU/NIC handles up to 8 concurrent flows, while with EP32 this rises to 24. With out-of-the-box DCQCN as congestion control, this causes severe throughput collapse.
As shown in Figure˜15(b), FAST outperforms the baseline by 1.75–7.88. Notably, FAST and RCCL exhibit opposite scaling with : increasing improves FAST by enlarging flows and amortizing staging overhead, but degrades RCCL due to increased flow collisions and congestion.
@@ 5.3 Scheduling Overhead
Scheduling runtime. As shown in Figure˜16, FAST scales to 320 GPUs with only 77 ms of overhead—faster than SyCCL, the fastest solver-based scheduler, which already takes 3.6s at just 16 GPUs. Earlier solver-based methods generally fail to scale beyond 64 GPUs, rendering them unusable for moderate expert-parallelism levels such as EP96 and EP128.
@@
Memory overhead. FAST also requires additional memory for temporary buffers that hold rebalanced or redistributed data. Under random workloads, this overhead is about of the original alltoallv buffer size. In practice, the impact is minimal: typical alltoallv buffers are under 1 GB (e.g., in DeepEP [deepep]), so the extra cost is less than 300 MB. On modern GPUs such as the NVIDIA H200 [h200] with 141 GB of memory, this represents under 0.22% of total capacity—an acceptable tradeoff for the performance gains.
5.4 Scaling and Bandwidth Sensitivity
@@
We use simulation to evaluate FAST beyond the limits of our physical testbeds, exploring both larger scales and different scale-up/scale-out bandwidth configurations. The simulator follows the analytical framework widely used in prior work such as TE-CCL and TACCL [taccl, teccl]: given a schedule with a sequence of transfer steps (each with a defined size), the completion time is computed by summing per-step costs. Each cost consists of a fixed link wake-up delay plus the transmission time ().
@@
We focus on scenarios that solver-based schedulers cannot scale to and therefore exclude them from comparison. Accordingly, we compare FAST against SpreadOut and an optimal bandwidth bound, which assumes infinitely fast scale-up links so that intra-server transfers are instantaneous. Under this bound, scale-out is the only bottleneck, and the optimal time is defined by the maximum balanced sender or receiver load divided by the scale-out bandwidth.
@@
Performance at larger scale. We first scale the number of GPUs in alltoallv, with each GPU pair transmitting 50 MB on average in a random workload, simulated in a 400 Gbps scale-out network and a 450 GBps scale-up network (H200). As shown in Figure˜17(a), FAST stays within 5% of optimal when scheduling time is excluded (“FAST raw”). Including scheduling time, the gap widens to 10% at larger scales, since scheduling cost grows faster than workload completion time (which increases only linearly with GPU count). We leave on-the-fly scheduling at extreme scale as future work. By contrast, SpreadOut achieves only about half of FAST ’s throughput.
@@
Performance under varying scale-up/scale-out ratios. We also evaluate FAST under different scale-up/scale-out bandwidth ratios on a 32-GPU setup. As shown in Figure˜17(b), normalized bandwidth is reported relative to scale-out capacity, which can exceed 1 since about 25% of traffic is intra-server, giving an upper bound near 1.25. Performance improves as the scale-up/scale-out ratio increases, as faster scale-up links further reduce balancing and redistribution overhead. These results indicate that FAST sustains high efficiency on modern GPUs with advanced scale-up interconnects.
6 Discussion
Other collectives. FAST is intentionally specialized for alltoallv, where skewness and dynamism fundamentally limit existing approaches, and is not designed as a general-purpose collective scheduler. For balanced collectives such as All-Reduce and All-Gather, communication patterns are static with uniform flow sizes, and near-optimal schedules can typically be achieved using existing library or solver-based implementations; a dynamic, traffic-aware scheduler provides little additional benefit. In practice, modern libraries (e.g., NCCL) already select among multiple collective algorithms at runtime, and FAST naturally fits this model as a specialized implementation for skewed alltoallv workloads.
Other topologies. FAST targets two-tier GPU cluster topologies that combine scale-up and scale-out fabrics, which are increasingly common in modern deployments [meta, junipter_network] with HGX-style platforms [hgx]. The core insight of FAST is to reshape traffic to reduce skew before it reaches slower network tiers, thereby simplifying the workload imposed on bottleneck fabric. While our current implementation focuses on two tiers, this principle naturally extends to multi-tier networks.
Hybrid parallelism. Our evaluation focuses on pure expert parallelism (EP). In hybrid deployments combining EP with tensor or pipeline parallelism, FAST can be applied directly when collectives execute in disjoint time windows. Coordinating FAST with concurrent collectives that share networks introduces additional complexity and is left for future work.
Persistence of scale-up vs. scale-out bandwidth gaps. We expect the bandwidth gap between scale-up and scale-out fabrics to persist, as scale-up interconnects benefit from short distances and tight hardware–software co-design, while scale-out networks must span longer distances and support backward compatibility and multi-tenancy. Exploiting scale-up bandwidth will therefore remain an important opportunity.
Interaction with congestion control. FAST operates at the collective layer and is orthogonal to transport-layer congestion control. Even under ideal congestion control, FAST can provide additional benefits by reducing traffic skew and improving NIC utilization, preventing communication from being bottlenecked by a small number of heavily loaded NICs.